### 24 Apr 2026 v1

# Effect of kinetic energy operator on double heterostructures spectra

Abstract
We report on the effect of the kinetic energy operator ambiguity on the energy spectra of various double heterostructures when the mass of the charge carriers, subjected to a potential, depends on position. The spectra are calculated using two complementary techniques. In the first case, which is not always possible, the energy spectrum satisfies a transcendental equation, which is obtained through an analytical process. While in the second, which can be used practically in any double heterostructure (DH), the energy spectra correspond to the poles of the reflection coefficient of the heterostructure. The spectra thus obtained complement those already reported in a previous reference, since we now include the spectra of another kinetic energy operator (KEO). Finally, we show how these methods allow us to critically analyze some statements that have been made about the relevance of some kinetic energy operators in some simpler heterostructures.
Keywords— Double heterostructure, effective position-depend mass, von Ross’s operator, boundary conditions, kinetic energy operator, energy spectra, reflection and transmission coefficients.
1 Introduction
When considering heterostructures in condensed matter physics, the low-energy excitations of charge carriers are described by an effective Schrödinger equation, where the KEO is an ambiguous quantity. This ambiguity is due to the fact that the mass operator $m=M_{0}m(z)$, no longer commutes with the momentum operator $\hat{p}=-i\hbar d/dz$, leaving the called von Ross KEO [34]
$T_{\textrm{vR}}(\alpha,\gamma)=\frac{1}{4}(m^{\alpha}\hat{p}m^{\beta}\hat{p}m^{\gamma}+m^{\gamma}\hat{p}m^{\beta}\hat{p}m^{\alpha}),$ (1)
without a defined ordering, because to guarantee its hermiticity, $\alpha+\beta+\gamma=-1$. This problem of the ambiguity of the von Roos KEO has been widely discussed in Refs. [31, 6, 33, 24, 25, 9, 16, 2, 14, 29, 11, 28, 21, 32, 17, 12, 13, 20, 5, 26], from different perspectives. In this work, we address some approaches that, despite having been replicated in some references, must be treated critically. For example, Ref. [31] concludes that energy eigenvalues of a particle jumping in mass within a finite one-dimensional box, diverge “unless $\alpha=\gamma$ and $\beta=-1$ is the only universal set of operator ordering parameters” with unique results. On the other hand, in the Ref. [6] some KEOs have been discarded. The main idea to do this is “to suppose that, once one have found the ordering without ambiguity for a given potential or class of potentials, that ordering should be extended to remaining physical potentials” and it warns that for an unconfined particle with position-depend mass its energy spectrum is “not well defined” for some of the orderings where $\alpha\neq\gamma$ and $\alpha=\gamma$.
For the purposes of this work, a graded heterostructure (GH) is characterized by the potential $V(z)$ and mass $m(z)$ distributions that change gradually and continuously throughout the structure. A double heterostructure (DH) is defined with mass profiles $m_{\textrm{in}}(z)$ and potential $V_{\textrm{in}}(z)$ graded in the intermediate region $z_{0}<z<z_{1}$, but they are constants in other regions,
$V(z)=\begin{cases}V_{0}\\ V_{\textrm{in}}(z)\\ V_{2}\end{cases},\qquad m(z)=\begin{cases}m_{0}&;\qquad z\leq z_{0},\\ m_{\textrm{in}}(z)&;\qquad z_{0}<z<z_{1},\\ m_{2}&;\qquad z_{1}\leq z.\end{cases}$ (2)
In Ref. [33], following two different but complementary strategies the effect of KEO ambiguity on double heterostructures spectra had been estimated for KEOs with $\alpha=\gamma$. In both procedures, the corresponding boundary conditions for the wave function of the charge carriers at points where there is a finite abrupt change in mass must be taken into account. That is, at the points where there is an abrupt finite change of mass (and / or potential) defined by the Heaviside function, $H$, as
$m(z)=m_{k}H(z_{k}-z)+m_{k+1}H(z-z_{k}),$ (3)
and at the points $z_{0,1}$ of the DH (2) when $m_{\textrm{in}}(z_{0,1})\neq m_{0,2}$.
In this Ref. [33], only the ambiguities $\alpha=\gamma=0$ of BenDaniel-Duke (BD-D) [3] and $\alpha=\gamma=-1/2$ of Zhu-Kroemer (Z-K) [36] were considered without taking into account choices such as $\beta=\gamma=-1/2$ of Li-Kuhn (L-K) [22] where $\alpha\neq\gamma$. Other examples of orderings where $\alpha\neq\gamma$ are given by Gora and Williams (G-W) [15] and the ordering that gives rise to the operator (L) [23]
$T_{L}=\frac{1}{6}\left(m^{-1}\hat{p}^{2}+\hat{p}m^{-1}\hat{p}+\hat{p}^{2}m^{-1}\right),$ (4)
which has the ambiguity defined by conditions $\alpha+\gamma=-\frac{2}{3}$, $\alpha\gamma=0$ and therefore is a particular case of von Ross KEO with $\alpha\neq\gamma$. For this KEO, $T_{L}$, the boundary conditions for the wave function $\psi(z)$ of charge carriers at points $z_{k}$ within the heterostructure, are [23]
$\psi(z_{k}^{-})=\frac{2m_{k}+m_{k+1}}{m_{k}+2m_{k+1}}\psi(z_{k}^{+});\qquad\psi^{\prime}(z_{k}^{-})=\frac{5m_{k}+m_{k+1}}{m_{k}+5m_{k+1}}\psi^{\prime}(z_{k}^{+}).$ (5)
Despite the warning in Ref. [31], the question arises: to which energy spectrum does a forbidden ordering lead when using any of the strategies of Ref. [33] for the DHs in this work?.
This paper is organized as follows. Section 2 briefly analyzes the ambiguity that leads to the KEO (4) and poses the Schrödinger equation to be solved for DH (2). Section 3 generalizes the strategies followed in Ref. [33], to now include the DHs energy spectra with orderings $\alpha\neq\gamma$. In Section 4, the article provides new energy spectra for the DHs of Ref. [33]. These spectra belong to a case of ambiguity that leads to operator (4), where, as we have said $\alpha\neq\gamma$ and the conditions (5) must be used. This section also includes two DHs, made with the distributions of the GHs of Ref. [6]. Additionally, we make some observations about this reference. Finally in Section 5, based on the energy spectra of the DHs obtained with three different KEO ambiguities, we give our conclusions.
2 Schrödinger equation for position-dependent effective masses
Before setting up the Schrödinger equation for the DH (2), let us observe the connection existing between the operators given by Eqs. (1) and (4), which can be rewritten as follows [^1] In the following we adopt units $\hbar^{2}(2M_{0})^{-1}=1$ and $m^{\prime}(z)$ denotes the derivative of $m(z)$ with respect to $z$.:
$T_{\textrm{vR}}(\alpha,\gamma)=-\frac{d}{dz}\frac{1}{m(z)}\frac{d}{dz}+V_{\alpha\gamma}(z);\qquad V_{\alpha\gamma}(z)=-\frac{1}{2}\left(\nu\frac{m^{\prime\prime}(z)}{m(z)^{2}}-\eta\frac{m^{\prime}(z)^{2}}{m(z)^{3}}\right),$ (6)
$\nu=\alpha+\gamma,\qquad\eta=2(\alpha+\gamma+\alpha\gamma)$, and
$T_{\textrm{L}}=-\frac{d}{dz}\frac{1}{m(z)}\frac{d}{dz}+V_{L}(z);\qquad V_{L}(z)=\frac{1}{3}\left(\frac{m^{\prime\prime}(z)}{m(z)^{2}}-2\frac{m^{\prime}(z)^{2}}{m(z)^{3}}\right),$ (7)
respectively. We observe that the operator $T_{\textrm{L}}$ is a particular case of $T_{\textrm{vR}}(\alpha,\gamma)$ if
$\nu=\alpha+\gamma=-\frac{2}{3};\qquad\eta=2(\alpha+\gamma+\alpha\gamma)=-\frac{4}{3}.$ (8)
That is, $\alpha$ and $\gamma$ must satisfy $\alpha+\gamma=-\frac{2}{3}$ and $\alpha\gamma=0$ (i.e. $\alpha\neq\gamma$). This conclusion differs from Ref. [23] which states that the $T_{L}$ operator is not a von Roos KEO.
Thus, the Schrödinger equation
$\left(-\frac{d}{dz}\frac{1}{m(z)}\frac{d}{dz}+V_{\alpha\gamma}(z)+V(z)-E\right)\psi=0,$ (9)
allow us to obtain the quantum states of the charge carrier of mass $m(z)$. As far as the energy spectrum is concerned $E$, the boundary conditions of $\psi$, are relevant. In the DH intermediate region must consider the boundary conditions at the points $z_{k}$ where it is fulfilled (3). In Ref. [33], where the von Roos KEO was considered with $\alpha=\gamma$, the boundary conditions for the wave function $\psi$ and its derivative at these points are given by [23]
$\psi(z_{k}^{-})=\frac{m_{k+1}^{\alpha+1}m_{k}^{\alpha+\beta+1}+m_{k}}{m_{k}^{\alpha+1}m_{k+1}^{\alpha+\beta+1}+m_{k+1}}\psi(z_{k}^{+});\qquad\psi^{\prime}(z_{k}^{-})=\frac{m_{k}^{\alpha+1}m_{k+1}^{\alpha+\beta+1}+m_{k}}{m_{k+1}^{\alpha+1}m_{k}^{\alpha+\beta+1}+m_{k+1}}\psi^{\prime}(z_{k}^{+}).$ (10)
On the other hand, the corresponding boundary conditions used for the KEO (7) are the Eqs. (5). In fact, Eqs. (5) and (10) explicitly show that each ordering of KEO, provides particular boundary conditions to the wave function at points $z_{k}$ which, as we have said, causes an effect on the energy spectrum $E$.
A possible way [^2] A variation of the method is outlined in Appendix B, applied to the DHs specified in Tables 8 and 9. to solve Eq. (9) is through its associated Schrödinger equation with unit mass, to see it we define
$\psi(z)=m(z)^{1/4}\phi,\qquad\rho=\int\sqrt{m(z)}dz,$ (11)
in said equation. The function $m(z)$ must allow the inverse transformation $z(\rho)$. The $\phi$ function satisfies the isospectral Schrödinger equation to the Eq. (9) with constant mass
$-\frac{d^{2}\phi}{d\rho^{2}}+(\widetilde{V}(\rho)-E)\phi=0,$ (12)
and potential
$\widetilde{V}(\rho)=V(z)+\frac{1}{2}\left(\eta+\frac{7}{8}\right)\frac{m^{\prime}(z)^{2}}{m(z)^{3}}-\frac{1}{2}\left(\nu+\frac{1}{2}\right)\frac{m^{\prime\prime}(z)}{m(z)^{2}}.$ (13)
Then, for appropriate $V(z)$ and $m(z)$, the associated Eq. (12) could be identified with a simpler or already known one. When this is the case, $E$ could be obtained directly and $\psi$ from $\phi$ and $m$ according to Eq. (11), where, in the cases of smooth profiles, the ‘generals’ boundary conditions must be taken into account as we will see in Section 3.1. This idea is one of the techniques already applied in Ref. [33], for two DHs that we again report in Secs. 4.1 and 4.2. The same idea has also been applied in Ref. [6] in two GHs to discard KEOs. In this work, each of these two GHs, is trapped in a DH (Eq. (2)), as described in Secs. 4.4, and 4.5. Given that from our DH models we can return to the GH model, we critically discuss the procedure followed in Ref [6] to discard KEOs.
3 Two strategies for calculating energy spectra in DHs
The DH model we have adopted (2) can be seen from two complementary points of view. In the first, when the Eq. (6) is taken into account, the Schrödinger equation (9)
$\left(-\frac{d}{dz}\frac{1}{m_{\textrm{in}}(z)}\frac{d}{dz}-\frac{1}{2}\left(\nu\frac{m_{\textrm{in}}^{\prime\prime}(z)}{m_{\textrm{in}}(z)^{2}}-\eta\frac{m_{\textrm{in}}^{\prime}(z)^{2}}{m_{\textrm{in}}(z)^{3}}\right)+V_{\textrm{in}}(z)-E\right)\psi_{\textrm{in}}=0,$ (14)
must be resolved by the wave function of carries. In this equation, the spatial distributions involved may or may not be continuous at the points $z_{k}$ of heterojunction. In practice, this equation can be solved analytically only for some potential and mass distributions.
In the second point of view, practically any pair of distributions $V(z)$ and $m(z)$ not necessarily confined, can be approximated as a succession of abrupt finite steps in potential and mass
[…]
${z_{\textrm{in}}}_{j}$; $j=0,1,\dots,n$ (${z_{\textrm{in}}}_{0}=z_{0},{z_{\textrm{in}}}_{n}=z_{1}$) represents a point at the intermediate region, where there is a finite abrupt change of potential and mass. Since charge carriers with position-dependent effective mass $m(z)$ within the DH are subject to the potential $V(z)$, their probability amplitude must be a solution of the simple Schrödinger equation
$-\frac{1}{m_{\textrm{in}j}}\frac{d^{2}}{dz^{2}}\psi_{\textrm{in}j}+(V_{\textrm{in}j}-E)\psi_{\textrm{in}j}=0,$ (16)
in each interval ${z_{\textrm{in}}}_{j-1},{z_{\textrm{in}}}_{j}]$. Evidently, the probability amplitude $\psi_{\textrm{in}j}$ of each interval must be joined with the neighboring amplitudes $\psi_{\textrm{in}j-1}$ and $\psi_{\textrm{in}j+1}$ , using the boundary conditions (5) and (10) that we have available for this work.
3.1 First strategy: ‘analytic’ method
According to Ref. [33], this is an analytical method to solve the Schrödinger equation (14). It can be used to verify the accuracy of the numerical methodology that will be presented in Section 3.2. We will achieve this by comparing the ‘analytical’ energy spectrum of a charge carrier with the poles of the reflection coefficient of the DH, as was previously done in Ref. [33] for cases $\alpha=\gamma$. In the inner region the particle obeys Eq. (14) so, the complete solution, in the case of bound states, can be written as
$\psi(z)=\begin{cases}\psi_{0}(z)=Re^{\eta_{0}z}&;\qquad z<z_{0},\\ \psi_{\textrm{in}}(z)=P\psi^{1}_{\textrm{in}}(z)+Q\psi^{2}_{\textrm{in}}(z)&;\qquad z_{0}\leq z\leq z_{1},\\ \psi_{2}(z)=Te^{-\eta_{2}z}&;\qquad z_{1}<z,\end{cases}$ (17)
where $\eta_{0,2}=\sqrt{m_{0,2}(V_{0,2}-E)}$. At points, $z_{0}$ and $z_{1}$ we must use boundary conditions (see. Ref. [33]) pending of the KEO we have chosen. Those conditions get us an equation system that we can write as
$\begin{cases}Re^{\eta_{0}z_{0}}=P\psi^{1}_{\textrm{in}}(z_{0})+Q\psi^{2}_{\textrm{in}}(z_{0})=-P\Phi_{11}-Q\Phi_{12},\\ Te^{-\eta_{2}z_{1}}=P\psi^{1}_{\textrm{in}}(z_{1})+Q\psi^{2}_{\textrm{in}}(z_{1})=-P\Phi_{21}-Q\Phi_{22}.\end{cases}$ (18)
This homogeneous system can be written in matrix form as
$X\left(\begin{array}[]{c}P\\ Q\end{array}\right):=\left(\begin{array}[]{cc}\psi^{1}_{\textrm{in}}(z_{0})+\Phi_{11}&\psi^{2}_{\textrm{in}}(z_{0})+\Phi_{12}\\ \psi^{1}_{\textrm{in}}(z_{1})+\Phi_{21}&\psi^{2}_{\textrm{in}}(z_{1})+\Phi_{22}\end{array}\right)\left(\begin{array}[]{c}P\\ Q\end{array}\right)=0.$ (25)
In general, these coefficients (functions) $\Phi_{ji}$, rely on the KEO’s choice. For the BD-D KEO we get
$\Phi_{1i}=-\frac{1}{\eta_{0}}\frac{m_{0}}{m_{\textrm{in}(z_{0})}}\left(\psi^{i}_{\textrm{in}}(z)\right)^{\prime}|_{z=z_{0}},\qquad\Phi_{2i}=\frac{1}{\eta_{2}}\frac{m_{2}}{m_{\textrm{in}(z_{1})}}\left(\psi^{i}_{\textrm{in}}(z)\right)^{\prime}|_{z=z_{1}};\qquad i=1,2.$ (26)
For the Z-K KEO
$\Phi_{1i}=-\frac{\sqrt{m_{\textrm{in}(z_{0})}}}{\eta_{0}}\left(\frac{\psi^{i}_{\textrm{in}}(z)}{\sqrt{m_{\textrm{in}(z)}}}\right)^{\prime}|_{z=z_{0}},\qquad\Phi_{2i}=\frac{\sqrt{m_{\textrm{in}(z_{1})}}}{\eta_{2}}\left(\frac{\psi^{i}_{\textrm{in}}(z)}{\sqrt{m_{\textrm{in}(z)}}}\right)^{\prime}|_{z=z_{1}};\qquad i=1,2.$ (27)
For a nontrivial solution of (25) the determinant of $X$ must be zero
$|X|=0.$ (28)
This is a transcendental equation that has been obtained through the analytical process described and allows us to calculate the energies of the bound states, as seen later in the DHs of Sections 4.1, 4.2, 4.4, and 4.5. The energy spectra of these analytical bound states are compared with the reflection coefficient poles which can be calculated using the multi-step methodology of Section 3.2.
When the DH of the Section 4.1 (4.2) is generalized through a parameter $\delta$ ($\tau$), as in Subsection 4.1.1 (4.2.1), here, it is no longer possible to obtain analytical spectra so, we estimate the bound states energy spectra only by the method described in Section 3.2.
On the other hand, from Eq. (18) we calculate
$P=-\frac{\Phi_{12}+\psi^{2}_{\textrm{in}}(z_{0})}{\Phi_{11}\psi^{2}_{\textrm{in}}(z_{0})-\Phi_{12}\psi^{1}_{\textrm{in}}(z_{0})}e^{\eta_{0}z_{0}}R,\qquad Q=\frac{\Phi_{11}+\psi^{1}_{\textrm{in}}(z_{0})}{\Phi_{11}\psi^{2}_{\textrm{in}}(z_{0})-\Phi_{12}\psi^{1}_{\textrm{in}}(z_{0})}e^{\eta_{0}z_{0}}R,\qquad T=(P\psi^{1}_{\textrm{in}}(z_{1})+Q\psi^{2}_{\textrm{in}}(z_{1}))e^{\eta_{2}z_{1}}.$ (29)
3.2 Second strategy: Multi-Step method
The reflection coefficient of the DH defined by the discrete distributions of potential and mass in the Eq. (15) is calculated. This is a numerical method exposed in Ref. [33] for the ordering $\alpha=\gamma$. The method solves the Schrödinger Eq. (16) in each interval ${z_{\textrm{in}}}_{j-1},{z_{\textrm{in}}}_{j}]$, taking into account the corresponding boundary conditions for the wave function $\psi(z)$ at each point where there is an abrupt finite change in mass and/or potential. This work including the case $\alpha\neq\gamma$ so generalizes the one published in Ref. [33].
In general, any spatial distribution of the potential $V(z)$ and of mass $m(z)$ can be seen as a succession of abrupt finite steps in the potential and mass. Some examples of this are the DHs in Section 4. Then, as we have already said, the values of the parameter pair $\alpha$ and $\gamma$ determine the appropriate boundary conditions for each interface. In this work, we have used the boundary conditions (5) and (10), with which the reflection amplitude $R$ of the charge carriers of the DH with discretized $V(z)$ and $m(z)$ distributions can be obtained recursively; first for one interface, then for two interfaces, and so on. The poles set of the reflection coefficient $R_{c}=|R|^{2}$, that can be obtained numerically, represents the energy spectrum of the DH.
For the above, we consider a potential function and position-dependent effective mass given by Eq. (15) and we calculate the reflection coefficient ($R$) for an arbitrary integer $n$; which must be large enough for the case of continuous distributions. For this purpose, we first consider the following, at a point $z_{j}$ where there is a change of mass (3) and potential, the boundary conditions can be written in general as (see Eqs. (5) and (10))
$\psi_{\textrm{in}j}(z_{\textrm{in}j}^{-})=\mu_{\textrm{in}j}\psi_{\textrm{in}j+1}(z_{\textrm{in}j}^{+});\qquad\psi^{\prime}_{\textrm{in}j}(z_{\textrm{in}j}^{-})=\rho_{\textrm{in}j}\psi^{\prime}_{\textrm{in}j+1}(z_{\textrm{in}j}^{+}),$ (30)
where $\mu_{\textrm{in}j}=\mu_{\textrm{in}j}(m_{\textrm{in}j},m_{\textrm{in}j+1})$ and $\rho_{\textrm{in}j}=\rho_{\textrm{in}j}(m_{\textrm{in}j},m_{\textrm{in}j+1})$ are functions of the ‘constant’ mass at the left ($m_{\textrm{in}j}$) and right ($m_{\textrm{in}j+1}$) of $z_{\textrm{in}j}$. $\psi_{\textrm{in}j}$ and $\psi_{\textrm{in}j+1}$ are the functions at the left and right of $z_{\textrm{in}j}$ respectively. The Schrödinger equation in $z_{\textrm{in}j-1},z_{\textrm{in}j}]$ is given by Eq. (16) where
$\psi_{\textrm{in}j}(z)=A_{\textrm{in}j}e^{ik_{\textrm{in}j}z}+B_{\textrm{in}j}e^{-ik_{\textrm{in}j}z},$ (31)
$k_{\textrm{in}j}=\sqrt{m_{\textrm{in}j}(E-V_{\textrm{in}j})}$. Applying the general boundary conditions (30) to the functions $\psi_{\textrm{in}j}$ and $\psi_{\textrm{in}j+1}$ at $z_{\textrm{in}j}$ we get the follow system of equations
$A_{\textrm{in}j}e^{ik_{\textrm{in}j}z_{\textrm{in}j}}+B_{\textrm{in}j}e^{-ik_{\textrm{in}j}z_{\textrm{in}j}}$ $=\mu_{\textrm{in}j}\left(A_{\textrm{in}j+1}e^{ik_{\textrm{in}j+1}z_{\textrm{in}j}}+B_{\textrm{in}j+1}e^{-ik_{\textrm{in}j+1}z_{\textrm{in}j}}\right),$ (32)
$A_{\textrm{in}j}e^{ik_{\textrm{in}j}z_{\textrm{in}j}}-B_{\textrm{in}j}e^{-ik_{\textrm{in}j}z_{\textrm{in}j}}$ $=\frac{k_{\textrm{in}j+1}}{k_{\textrm{in}j}}\rho_{\textrm{in}j}\left(A_{\textrm{in}j+1}e^{ik_{\textrm{in}j+1}z_{\textrm{in}j}}-B_{\textrm{in}j+1}e^{-ik_{\textrm{in}j+1}z_{\textrm{in}j}}\right).$ (33)
From Eqs. (32) and (33) we obtain
$\frac{B_{\textrm{in}j}}{A_{\textrm{in}j}}=\frac{r_{j,j+1}+\frac{B_{\textrm{in}j+1}}{A_{\textrm{in}j+1}}e^{-2ik_{\textrm{in}j+1}z_{\textrm{in}j}}}{1+r_{j,j+1}\frac{B_{\textrm{in}j+1}}{A_{\textrm{in}j+1}}e^{-2ik_{\textrm{in}j+1}z_{\textrm{in}j}}}e^{2ik_{\textrm{in}j}z_{\textrm{in}j}};\qquad r_{j,j+1}=\frac{k_{\textrm{in}j}\mu_{\textrm{in}j}-k_{\textrm{in}j+1}\rho_{\textrm{in}j}}{k_{\textrm{in}j}\mu_{j}+k_{\textrm{in}j+1}\rho_{\textrm{in}j}}.$ (34)
  • $R$ for a single interfaces ($n=1$)
    For a single interface, we have two functions $\psi_{\textrm{in}0}(z)=A_{\textrm{in}0}e^{ik_{\textrm{in}0}z}+B_{\textrm{in}0}e^{-ik_{\textrm{in}0}z}$ and $\psi_{\textrm{in}1}(z)=A_{\textrm{in}1}e^{ik_{\textrm{in}1}z}+B_{\textrm{in}1}e^{-ik_{\textrm{in}1}z}$ whose coefficients satisfy the relation (34) with $B_{\textrm{in}1}=0$ and $z_{\textrm{in}0}=0$, so
    $R=\frac{B_{\textrm{in}0}}{A_{\textrm{in}0}}=r_{01}.$ (35)
  • $R$ for two single interfaces ($n=2$)
    In this case, we have three functions $\psi_{\textrm{in}0}(z)=A_{\textrm{in}0}e^{ik_{\textrm{in}0}z}+B_{\textrm{in}0}e^{-ik_{\textrm{in}0}z}$, $\psi_{\textrm{in}1}(z)=A_{\textrm{in}1}e^{ik_{\textrm{in}1}z}+B_{\textrm{in}1}e^{-ik_{\textrm{in}1}z}$, and $\psi_{\textrm{in}2}(z)=A_{\textrm{in}2}e^{ik_{\textrm{in}2}z}$ ($B_{\textrm{in}2}=0$). Using Eq. (34) we have the relationships
    $\frac{B_{\textrm{in}0}}{A_{\textrm{in}0}}=\frac{r_{01}+\frac{B_{\textrm{in}1}}{A_{\textrm{in}1}}e^{-2ik_{\textrm{in}1}z_{\textrm{in}0}}}{1+r_{01}\frac{B_{\textrm{in}1}}{A_{\textrm{in}1}}e^{-2ik_{\textrm{in}1}z_{\textrm{in}0}}}e^{2ik_{\textrm{in}0}z_{\textrm{in}0}},\qquad\frac{B_{\textrm{in}1}}{A_{\textrm{in}1}}=\frac{r_{12}+\frac{B_{\textrm{in}2}}{A_{\textrm{in}2}}e^{-2ik_{\textrm{in}2}z_{\textrm{in}1}}}{1+r_{12}\frac{B_{\textrm{in}2}}{A_{\textrm{in}2}}e^{-2ik_{\textrm{in}2}z_{\textrm{in}1}}}e^{2ik_{\textrm{in}1}z_{\textrm{in}1}}=r_{12}e^{2ik_{\textrm{in}1}z_{\textrm{in}1}}.$ (36)
    Thus, defining $w_{\textrm{in}1}=z_{\textrm{in}1}-z_{\textrm{in}0}$ and $z_{\textrm{in}0}=0$, the reflection coefficient is
    $R=\frac{B_{\textrm{in}0}}{A_{\textrm{in}0}}=\frac{r_{01}+r_{12}e^{2ik_{\textrm{in}1}w_{\textrm{in}1}}}{1+r_{01}r_{12}e^{2ik_{\textrm{in}1}w_{\textrm{in}1}}}.$ (37)
  • $R$ for three single interfaces ($n=3$)
    Now, we consider that the space is divided by three interfaces, we will have four wave functions whose coefficients will satisfy the relations
    $\frac{B_{\textrm{in}0}}{A_{\textrm{in}0}}=\frac{r_{01}+\frac{B_{\textrm{in}1}}{A_{\textrm{in}1}}e^{-2ik_{\textrm{in}1}z_{\textrm{in}0}}}{1+r_{01}\frac{B_{\textrm{in}1}}{A_{\textrm{in}1}}e^{-2ik_{\textrm{in}1}z_{\textrm{in}0}}}e^{2ik_{\textrm{in}0}z_{\textrm{in}0}},\qquad\frac{B_{\textrm{in}1}}{A_{\textrm{in}1}}=\frac{r_{12}+\frac{B_{\textrm{in}2}}{A_{\textrm{in}2}}e^{-2ik_{\textrm{in}2}z_{\textrm{in}1}}}{1+r_{12}\frac{B_{\textrm{in}2}}{A_{\textrm{in}2}}e^{-2ik_{\textrm{in}2}z_{\textrm{in}1}}}e^{2ik_{\textrm{in}1}z_{\textrm{in}1}},\qquad\frac{B_{\textrm{in}2}}{A_{\textrm{in}2}}=r_{23}e^{2ik_{\textrm{in}2}z_{\textrm{in}2}}.$ (38)
    such that the reflection coefficient can be expressed in the following manner
    $R=r_{0123}=\frac{B_{\textrm{in}0}}{A_{\textrm{in}0}}=\frac{r_{01}+r_{123}e^{2ik_{\textrm{in}1}w_{\textrm{in}1}}}{1+r_{01}r_{123}e^{2ik_{\textrm{in}1}w_{\textrm{in}1}}},$ (39)
    where
    $r_{123}=\frac{r_{12}+r_{23}e^{2ik_{\textrm{in}2}w_{\textrm{in}2}}}{1+r_{12}r_{23}e^{2ik_{\textrm{in}2}w_{\textrm{in}2}}},$ (40)
    $w_{\textrm{in}2}=z_{\textrm{in}2}-z_{\textrm{in}1}$, $w_{\textrm{in}1}=z_{\textrm{in}1}-z_{\textrm{in}0}$, $B_{\textrm{in}3}=0$, and $z_{\textrm{in}0}=0$.
  • $R$ for $n$ interfaces
    The reflection amplitud of the system described by Eq. (15) is
    $R=r_{012\dots n}=\frac{B_{\textrm{in}0}}{A_{\textrm{in}0}}=\frac{r_{01}+r_{12\dots n}e^{2ik_{\textrm{in}1}w_{\textrm{in}1}}}{1+r_{01}r_{12\dots n}e^{2ik_{\textrm{in}1}w_{\textrm{in}1}}},$ (41)
    where
    $r_{12\dots n}$ $=\frac{r_{12}+r_{23\dots n}e^{2ik_{\textrm{in}2}w_{\textrm{in}2}}}{1+r_{12}r_{23\dots n}e^{2ik_{\textrm{in}2}w_{\textrm{in}2}}},$
    $r_{23\dots n}$ $=\frac{r_{23}+r_{34\dots n}e^{2ik_{\textrm{in}3}w_{\textrm{in}3}}}{1+r_{23}r_{34\dots n}e^{2ik_{\textrm{in}3}w_{\textrm{in}3}}},$
    $\vdots$
    $r_{n-1,n}$ $=\frac{k_{\textrm{in}n-1}\mu_{\textrm{in}n-1}-k_{\textrm{in}n}\rho_{\textrm{in}n-1}}{k_{\textrm{in}n-1}\mu_{\textrm{in}n-1}+k_{\textrm{in}n}\rho_{\textrm{in}n-1}}.$ (42)
    In this equation we have considered $z_{\textrm{in}0}=0$, $w_{\textrm{in}j}=z_{\textrm{in}j}-z_{\textrm{in}j-1}$, $\psi_{\textrm{in}n}(z)=A_{\textrm{in}n}e^{ik_{\textrm{in}n}z}$ $(B_{\textrm{in}n}=0)$, and $k_{\textrm{in}j}=\sqrt{m_{\textrm{in}j}(E-V_{\textrm{in}j})}$.
This reflection amplitud $R$ (i.e. reflection coefficient $R_{c}=|R|^{2}$) depends of the constants $\mu_{\textrm{in}j}$ and $\rho_{\textrm{in}j}$ through $r_{j,j+1}$. For example, if we used the KEO $T_{\textrm{vR}}$ (1) then (see Eq. (10))
$\mu_{\textrm{in}j}=\frac{m_{\textrm{in}j+1}^{\alpha+1}m_{\textrm{in}j}^{\alpha+\beta+1}+m_{\textrm{in}j}}{m_{\textrm{in}j}^{\alpha+1}m_{\textrm{in}j+1}^{\alpha+\beta+1}+m_{\textrm{in}j+1}};\qquad\rho_{\textrm{in}j}=\frac{m_{\textrm{in}j}^{\alpha+1}m_{\textrm{in}j+1}^{\alpha+\beta+1}+m_{\textrm{in}j}}{m_{\textrm{in}j+1}^{\alpha+1}m_{\textrm{in}j}^{\alpha+\beta+1}+m_{\textrm{in}j+1}}.$ (43)
In the case that we choose the values $\alpha=0$ and $\beta=-1$ ($\nu=\eta=0$), i.e. BenDaniel-Duke operator,
$\mu_{\textrm{in}j}=1;\qquad\rho_{\textrm{in}j}=\frac{m_{\textrm{in}j}}{m_{\textrm{in}j+1}}\qquad\Rightarrow\qquad r_{j,j+1}=\frac{\frac{k_{\textrm{in}j}}{m_{\textrm{in}j}}-\frac{k_{\textrm{in}j+1}}{m_{\textrm{in}j+1}}}{\frac{k_{\textrm{in}j}}{m_{\textrm{in}j}}+\frac{k_{\textrm{in}j+1}}{m_{\textrm{in}j+1}}}.$ (44)
On the other hand, for the values $\alpha=-1/2$ and $\beta=0$ ($\nu=-1$, $\eta=-3/2$), i.e. Zhu-Kroemer operator,
$\mu_{\textrm{in}j}=\rho_{\textrm{in}j}=\frac{m_{\textrm{in}j}^{1/2}}{m_{\textrm{in}j+1}^{1/2}}\qquad\Rightarrow\qquad r_{j,j+1}=\frac{k_{\textrm{in}j}-k_{\textrm{in}j+1}}{k_{\textrm{in}j}+k_{\textrm{in}j+1}}.$ (45)
If we used $\alpha=0$ and $\gamma=-2/3$ ($\nu=-2/3$, $\eta=-4/3$), i.e. $T_{L}$ operator, the coefficients $\mu_{\textrm{in}j}$ and $\rho_{\textrm{in}j}$ are given by Eq. (5)
$\mu_{\textrm{in}j}=\frac{2m_{\textrm{in}j}+m_{\textrm{in}j+1}}{m_{\textrm{in}j}+2m_{\textrm{in}j+1}};\qquad$ $\rho_{\textrm{in}j}=\frac{5m_{\textrm{in}j}+m_{\textrm{in}j+1}}{m_{\textrm{in}j}+5m_{\textrm{in}j+1}}$
$\qquad\Rightarrow\qquad r_{j,j+1}=\frac{\frac{k_{\textrm{in}j}}{(m_{\textrm{in}j}+2m_{\textrm{in}j+1})(5m_{\textrm{in}j}+m_{\textrm{in}j+1})}-\frac{k_{\textrm{in}j+1}}{(m_{\textrm{in}j}+5m_{\textrm{in}j+1})(2m_{\textrm{in}j}+m_{\textrm{in}j+1})}}{\frac{k_{\textrm{in}j}}{(m_{\textrm{in}j}+2m_{\textrm{in}j+1})(5m_{\textrm{in}j}+m_{\textrm{in}j+1})}+\frac{k_{\textrm{in}j+1}}{(m_{\textrm{in}j}+5m_{\textrm{in}j+1})(2m_{\textrm{in}j}+m_{\textrm{in}j+1})}}.$ (46)
So, regardless of the value of the coefficients $\mu_{\textrm{in}j}$ and $\rho_{\textrm{in}j}$ in the definition of $r_{j,j+1}$, the recursion form of the reflection coefficient $R$, remains. If $m_{\textrm{in}j}=m_{\textrm{in}j+1}$ then $\mu_{\textrm{in}j}=\rho_{\textrm{in}j}=1$ in all cases.
4 Double heterostructure models
Each Table in this section defines a DH. In them, three energy spectra are reported; the first two with $\alpha=\gamma$ and the third with $\alpha\neq\gamma$. The first of the columns indicates the calculation strategy used. We have highlighted the energy spectra obtained in this work for $\alpha\neq\gamma$. Appendix A is a complement to the DH defined in Tables 2 and 5, in which the analytical method strategy is applied. Tables 8 and 9 contain, in their intermediate region, the potential and mass distributions given in the Ref. [6].
4.1 Symmetric distributions of potential and mass
Table 2 defines the DH and shows its different energy spectra; the first row defines its distributions according to the notation of equation (2) and displays their graphs. The table presents the energy spectra for three different ambiguities; the first two correspond to the case $\alpha=\gamma$ and the third to a case $\alpha\neq\gamma$. The spectra with $\alpha=\gamma$ were taken from [33], and the spectra highlighted in the table corresponding $\alpha\neq\gamma$ are a result of this work. The spectra obtained through the solution of the transcendental equation (28) and those denoted by $\mathbf{E}_{n}$ are explained in the Appendix A.1.
4.1.1 Symmetric-Gaussian profiles
We now consider some modifications in the DH considered in the Table 2. Table 3 (4) defines a family of DHs and reports different energy spectra. The first row considers a single potential (mass) for different mass (potential) distributions, characterized through parameter $\delta$. This row also includes some graphs of the masses (potentials), and the effect of parameter $\delta$.
Because analytical solutions are difficult to find, we analyze these systems numerically, meaning we calculate the energy spectrum only using the poles of the reflection coefficient (41) for the different conditions of ambiguity. Tables 3 and 4 present the energy spectra where the first two rows correspond to two cases where $\alpha=\gamma$ and the third, to a case $\alpha\neq\gamma$. The spectra with $\alpha=\gamma$ were taken from [33], and the spectra corresponding to the case $\alpha\neq\gamma$ are reported in this work.
4.2 Morse-like distribution of potential and mass
Table 5 defines the DH and shows its different energy spectra; the first row defines its distributions according to the notation of equation (2) and displays their graphs. The table presents the energy spectra for three different ambiguities; the first two correspond to the case $\alpha=\gamma$ and the third to a case $\alpha\neq\gamma$. The spectra with $\alpha=\gamma$ were taken from [33], and the spectrum highlighted in the table corresponding to $\alpha\neq\gamma$ is a result of this work. The spectrum obtained through the solution of the transcendental Eq. (28) is explained in the Appendix A.2.
4.2.1 Truncated Morse potential with hyperbolic mass profile
We now consider some modifications in the DH considered in the Table 5. Table 6 defines a family of DHs and reports different energy spectra. The first row considers a single potential for different mass distributions, characterized through parameter $\tau$. This row also includes some graphs of the masses, highlighting the effect of parameter $\tau$.
Because analytical solutions are difficult to find, we analyze these systems numerically, meaning we calculate the energy spectrum only using the poles of the reflection coefficient (41) for the different conditions of ambiguity. Table 6 present the energy spectra where the first two correspond to the case $\alpha=\gamma$ and the third, to a case $\alpha\neq\gamma$. The spectra with $\alpha=\gamma$ were taken from [33], and the spectrum corresponding to the case $\alpha\neq\gamma$ is a result of this work.
4.3 Double parabolic quantum well
Table 7 shows different energy spectra of DH defined by
$V_{p}(z)=V_{0}\begin{cases}1\\ f(z)\\ g(z)\\ 1\end{cases},\qquad m_{p}(z)=\begin{cases}m_{0}&;z<b,\\ m_{1}+(m_{0}-m_{1})f(z)&;b\leq z<c,\\ m_{1}+(m_{0}-m_{1})q(z)&;c\leq z<d,\\ m_{0}&;d\leq z.\end{cases}$ (47)
where $f(z)=\frac{(z-c_{1})^{2}}{c_{2}^{2}}$, $g(z)=\frac{(z-c_{3})^{2}}{c_{4}^{2}}$; $c_{1}=\frac{c+a}{2},\qquad c_{2}=\frac{c-a}{2},\qquad c_{3}=\frac{d+c}{2},\qquad c_{4}=\frac{d-c}{2}$. First row in Table 7 displays their graphs. Similar to the cases where an analytical solution is not available, the energy spectrum was determined numerically through the poles of the reflection coefficient (41). The results for the cases $\alpha=\gamma$ were taken from [33], while the values for $\alpha\neq\gamma$ are newly obtained in the present work.
4.4 Exponential potential and mass profiles
In this section, we examined a modified profile studied in [6]. For our DH model (2) we define
$V(z)=\begin{cases}V_{0}=\lambda V_{2}\\ V_{\textrm{in}}(z)=V_{c}e^{cz}\\ V_{2}=V_{\textrm{in}}(z_{1})\end{cases},\qquad m(z)=\begin{cases}m_{0}=m_{\textrm{in}}(z_{0})&;\qquad z\leq z_{0},\\ m_{\textrm{in}}(z)=\mu_{0}e^{cz}&;\qquad z_{0}<z<z_{1},\\ m_{2}=m_{\textrm{in}}(z_{1})&;\qquad z_{1}\leq z.\end{cases}$ (48)
$V_{c},\mu_{0}>0$. It is noteworthy that $V_{0}=\lambda V_{2}$; $\lambda\in\mathbb{R}$. The potential distribution $V(z)$ is illustrated in Fig. 1. The corresponding Eq. (12) in the inner region $z_{0}<z<z_{1}$ is
$-\frac{d^{2}\phi}{d\rho^{2}}+\widetilde{V}(\rho)\phi=E\phi;\qquad\rho=\frac{2\sqrt{\mu_{0}}}{c}e^{cz/2}>0,$ (49)
$g=\frac{(3+8\eta-8\nu)}{2}$, $\omega^{2}=\frac{c^{2}V_{c}}{\mu_{0}}$, and
$\widetilde{V}(\rho)=\frac{1}{4}\omega^{2}\rho^{2}+\frac{1}{2}\frac{g}{\rho^{2}}.$ (50)
We would like to make the following comments on this potential to clarify some conclusions that were made in Ref. [6]. The parameters $V_{c}$, $c$ and $\mu_{0}$ of the distributions (48) define the positive parameter $\omega$. Independently, the ambiguity parameters $\alpha$ and $\gamma$ determine $g$. When $g$ is greater or less than zero, the potential takes the forms described in Table 1. In principle, both potentials are of completely different nature:
  • The term $g/\rho^{2}$ is a repulsive (attractive) potential barrier with a positive (negative) singularity at the origin only when its coefficient $g\geq 0$ ($g<0$). Nevertheless, this term retains the isochronous property that characterizes the harmonic oscillator when $g\geq-1/2$. This restriction includes potentials with $1/2\leq g<0$, where the intensity of the centripetal coupling is not strong enough to overcome the harmonic potential barrier. The energy spectrum of this isotonic potential is [4, 35, Cariñena_2008, 18, 30].
    $E=\omega\left(2n+1+d\right);\qquad n=0,1,\dots$ (51)
    where
    $d=\frac{1}{2}\sqrt{1+2g};\qquad g\geq-1/2.$ (52)
    which, as we see, is discrete and positive. In fact, $d$ is basically a constant for the factorization of the Hamiltonian in (49) and which according to the references [7, 8, 27] is real and positive. It is straightforward to verify that the operators of G-W, and $T_{L}$ lead to negative $g$ values less than 1/2-1/2. Therefore, the graded or double heterostructure that uses these KEOs does not have the energy spectrum (51). For this reason, conclusions about these KEOs used for (48) cannot be drawn from this energy spectrum.
  • On the other hand, when g<1/2g<-1/2, the isotonic term ceases to be so. The potential becomes more attractive by modifying the singularity at the origin in such a way that it must be addressed with special methods for its regularization, with which diverse energy spectra are obtained [Fernández_2021, 19]. These will be the spectra that must be related to the energy spectra of the heterostructure (48) when the G-W and $T_{L}$ operators are considered. This matter will be detailed in another work.
The Table 1, highlights that the graph of the isotonic potential with $1/2\leq g<0$ qualitatively coincides with the graph of the non-isotonic potential g<1/2g<-1/2 but their energy spectra are completely different. The only KEOs with proper name, which define a $g\geq-1/2$, are the Z-K, L-K (g=1/2g=-1/2) and BD-D ($g=3/2$) operators, and therefore the other ambiguities in (51) should not be used.
Figure 1: Potential $V(z)$ (gray line) of DH (48) and its eigenfunctions; $\psi^{1}_{\textrm{in}}(z)$ and $\psi^{2}_{\textrm{in}}(z)$ are given by Eq. (54). The energy spectrum $\{E_{0},E_{1},E_{2},E_{3}\}=\{4.63268,10.0389,15.223,20.076\}$ how was obtained using Eq. (28) correspond to the Z-K conditions accordingly with Table 8. Note that the obtained spectrum is discrete and positive, just as the isotonic spectrum (51) is, consistent with Table 1; $V_{c}=3$, $\mu_{0}=1/2$, $c=1$, $\lambda=1$, $z_{0}=-2$, $z_{1}=2$.
$\widetilde{V}(\rho)$ Energy spectrum $g$ value for each KEO
g<1/2g<-1/2 non-Isotonic 5/2-5/2 (G-W), 7/6-7/6 (L)
$1/2\leq g<0$ Isotonic Eq. (51) 1/2-1/2 (Z-K and L-K)
$0\leq g$ Isotonic Eq. (51) $3/2$ (BD-D)
Table 1: Potential $\widetilde{V}(\rho)$ (50) for each of the KEO ambiguities. The potential with $1/2\leq g<0$ is isotonic because it has the energy spectrum (51), although, qualitatively, its graph corresponds to that of the non-isotonic potential (g<1/2g<-1/2).
We find the solutions of (49) using the ansatz $\phi=y^{-1/4}F(y)$, where $F(y)$ satisfy the Whittaker equation
$\frac{d^{2}F}{dy^{2}}+\left(-\frac{1}{4}+\frac{\kappa}{y}+\frac{1/4-\mu^{2}}{y^{2}}\right)F=0;\qquad y=\frac{\omega}{2}\rho^{2},$ (53)
$\kappa=E/(2\omega)$, $\mu^{2}=\frac{1}{16}(1+2g)=\frac{1}{4}(1+2\eta-2\nu)$. Taking into account the Whittaker solutions of first and second kind $M_{\kappa,\mu}(y)$, $W_{\kappa,\mu}(y)$ of (53) respectively, we find (11)
$\psi_{\text{in}}^{1}(z)=M_{\kappa,\mu}(y);\qquad\psi_{\text{in}}^{2}(z)=W_{\kappa,\mu}(y);\qquad y=\frac{2\sqrt{V_{c}\mu_{0}}}{c}e^{cz}.$ (54)
In Figure 1, we present the first four non-normalized eigenfunctions and their corresponding energies, derived from Eq. (28) and Z-K boundary condition. The comprehensive energy spectrum for the KEOs investigated in this study is depicted in Table 8. In the cases where both calculation techniques were used, we observed consistent energy spectra. An important remark is that the numerical energy spectra depends on the parameter value $V_{0,1}$ as well as the values $z_{0}$ and $z_{1}$. These energy values are real for all KEOs of this work. Choosing some appropriate values of the parameters $z_{0}=-5$, $z_{1}=4$, and $\lambda=1$, the numerical energy values agree with Eq. (51) for the BD-D condition, that is to say, $g=3/2\Rightarrow E:=E^{BD-D}=w(2n+2)$.
We have verified that when the heterostructure ceases to be double, that is $|z_{0,1}|\rightarrow\infty$ and $V_{0}\rightarrow\infty$, and because of this $0<\rho<\infty$ (Eq. (49)), the energy spectrum of the system (48) that matches the spectrum of the isotonic oscillator (51) is that of the ambiguity defined by the KEO of BD-D.
4.5 Singular potential and parabolic mass
Finally, we examined the second profile studied in [6] given by
$m_{\textrm{in}}(z)=cz^{2};\qquad V_{\textrm{in}}(z)=\frac{A}{c}\left(\frac{1}{z^{4}}+\frac{B}{z^{2}}\right),\qquad z,c,A,-B>0.$ (55)
We defined $V_{0,2}=V_{\textrm{in}}(z_{0,1})$, $m_{0,2}=m_{\textrm{in}}(z_{0,1})$ for our model (2). In the inner region we get the associated equation (12)
$-\frac{d^{2}\phi}{d\rho^{2}}+\left(\frac{F}{\rho^{2}}+\frac{G}{\rho}-E\right)\phi=0;\qquad\rho=\frac{\sqrt{c}}{2}z^{2},$ (56)
$F=(5+8\eta-4\nu+4A)/16$, $G=AB/(2\sqrt{c})$, and E=|E|E=-|E|. The conditions $F>0$ and $G<0$ must be satisfied to generate a quantum well as the Fig. 2(b). The first of these conditions links the ambiguity parameters with the potential parameter $A$, that is,
$A>-\frac{5}{4}-(2\eta-\nu).$ (57)
This represents a necessary condition that must be considered in advance, before drawing any conclusions based on the energy spectrum of this well. Eq. (56) can be rewritten as the Whittaker equation
$\frac{d^{2}\phi}{dy^{2}}+\left(-\frac{1}{4}+\frac{\kappa}{y}+\frac{1/4-\mu^{2}}{y^{2}}\right)\phi=0,$ (58)
where $y=2\sqrt{|E|}\rho$, $\kappa=-G/(2\sqrt{|E|})$, $\mu^{2}=\frac{1}{4}+F$. Whose solutions are the Whittaker functions of the first and second kind $M_{\kappa,\mu}(y)$, $W_{\kappa,\mu}(y)$ respectively. So, the exact solutions (11) in the inner region are
$\psi_{\text{in}}^{1}(z)=c^{1/4}\sqrt{z}M_{\kappa,\mu}(y);\qquad\psi_{\text{in}}^{2}(z)=c^{1/4}\sqrt{z}W_{\kappa,\mu}(y).$ (59)
The boundary conditions consistent with the potential well are $\psi(z=0,\infty)=0$. In Fig. 2(a), we plot the complete eigenfunctions ($\infty<z<\infty$) of our model (2) for some parameter values.
(a) Potential $V(z)$ of DH (55). Conditions A,B>0A,-B>0 must be satisfied to generate a quantum well.
(b) Potential $\widetilde{V}(\rho)$ of Hamiltonian (56). Conditions F,G>0F,-G>0 must be satisfied to generate a quantum well.
Figure 2: a) First three eigenfunctions (17) of the system (55), $\psi_{\text{in}}^{1}(z)$, $\psi_{\text{in}}^{1}(z)$ are given by Eqs. (59) and its energy values $\{E_{0},E_{1},E_{2}\}=\{-16.058849,-4.948710,-2.370368\}$ correspond to Z-K condition; $A=2$, B=10B=-10, $c=1$, $z_{0}=0.1$, $z_{1}=4$.
We have verified that when the heterostructure ceases to be double and is reduced to the GH of the Ref. [6], that is, in the limit $z_{0}\rightarrow 0$ (i.e. $V_{0}\rightarrow\infty$) and $z_{1}\rightarrow\infty$ $\Rightarrow 0<\rho<\infty$, the analytical energy spectrum of the potential (56) is (see Eq. (B.8) in Appendix B)
$E=-\frac{(AB)^{2}}{4c\left(2n+1+\frac{1}{2}\sqrt{1+2g}\right)^{2}};\qquad g=2(2+2\eta-\nu)+2A,\qquad n=0,1,\dots$ (60)
Evidently, the sufficient condition to guarantee the reality of the energy spectrum is
$g>-\frac{1}{2}.$ (61)
Note that if the ambiguity parameters meet the sufficient condition (61) but not the necessary condition (57), the energy spectrum (60) ceases to make sense. So, for all conditions, we can always choose an appropriate value for $A>-\frac{5}{4}-(2\eta-\nu)>-\frac{9}{4}-(2\eta-\nu)$ such that the energy spectrum (60) is real for all ambiguity
${\mathbf{E}_{n}}^{BD-D}=-\frac{(AB)^{2}}{4c\left(2n+1+\sqrt{9/4+A}\right)^{2}},\qquad{\mathbf{E}_{n}}^{Z-K}=-\frac{(AB)^{2}}{4c\left(2n+1+\sqrt{1/4+A}\right)^{2}}.$ (62)
We can observe that the parameter values associated with the $T_{L}$ KEO ($\nu=-2/3,\eta=-4/3$) lead us to the energy spectrum ${\mathbf{E}_{n}}^{T_{L}}={\mathbf{E}_{n}}^{Z-K}$ in Eq. (60). This observation is only applicable to the GH of the Ref. [6] and not to the case of DH (55), where the boundary conditions at points $z_{0}$ and $z_{1}$ for each of these KEO are different. Therefore, the energy spectra for these KEOs are different, as we can see in Table 9.
5 Conclusions
In this work, the strategies from Ref. [33] are generalized to include energy spectra where the KEOs have parameters such that $\alpha\neq\gamma$. That is, we have added a case of ambiguity that had not been considered in that reference. There are two reasons why we consider that the multi-step method described in Section 3.2 is an appropriate procedure for addressing ambiguities in DHs: 1) We have validated both strategies in [33], showing that the energy spectra of the double heterostructures that admit both procedures are practically identical, and 2) for all the DHs reported here, the numerical multi-step method is sensitive to the different ambiguities considered and that the orderings $\alpha\neq\gamma$ report finite and well-defined values close to those of the ordering $\alpha=\gamma$. Under these circumstances, the tables that define each of our double heterostructures can be used to select the kinetic energy operator that leads to the lowest energy spectrum. This selection criterion can be easily implemented in other types of heterostructures. In this way, we observe that the choice of the Z-K operator generally leads to the lowest energy spectra in the double heterostructures defined in Tables 2, 3, 4, 5, 6, and 9. It is the BD-D and L operators that determine the lowest energy spectra of the double heterostructures defined in Tables 7 and 8, respectively.
We have used the ordering of the kinetic energy operator that has been proposed in reference [23] and have also shown that von Ross’s KEO includes it. We have not found double heterostructures with ambiguities that prevent obtaining well-defined energy spectra, contrary to what has been reported in other references. For example, Ref. [31] establishes $\alpha=\gamma=0$ as the universally consistent ordering for KEO in effective mass theory, ensuring physically meaningful spectra for particle jumping in mass within a finite one-dimensional box. According to our results, we consider that the conclusions made in this reference are no longer valid for the type double heterostructures that we have adopted in this work. At the same time, we have critically discussed some of the arguments that were used for the conclusions made in the Ref. [6].
Acknowlegments
RVT acknowledges Secretaría de Ciencia, Humanidades, Tecnología e Innovación (SECIHTI - México) support under the grant FORDECYT-PRONACES/61533/2020. JGR acknowledges IPN project SIP 20260578 and SECIHTI. JA acknowledges IPN project SIP20260576 and SECIHTI.
Appendix A
In this appendix, the analytical method of Section 3.1 is applied to the double heterostructures of Sections 4.1 and 4.2. The energy spectra presented in Tables 2 and 5 are explained here.
A.1 Complement of the DH defined in Table 2.
We study the symmetric DH model with potential and mass distributions in the inner region $z_{0}<z<z_{1}$ in Eq. (2)
$V_{\textrm{in}}(z)=-\frac{\mu^{2}}{1+z^{2}},\qquad m_{\textrm{in}}(z)=\frac{\sigma^{2}}{1+z^{2}},$ (A.1)
$\mu$, $\sigma$ are non-zero real parameters. We define $z_{0}<0$, $z_{1}=|z_{0}|$, $V_{0}=V_{2}=V_{\textrm{in}}(z_{0})$, and $m_{0}=m_{2}=m_{\textrm{in}}(z_{0})$, so that $V(z)$ and $m(z)$ are continuous as is shown in figure inserted in Table 2. The associated Schrödinger equation (12) for this system can be written as
$\frac{d^{2}\phi}{d\rho^{2}}+\left(\kappa^{2}+\frac{\lambda(\lambda-1)}{\sigma^{2}}\frac{1}{\cosh^{2}{\frac{\rho}{\sigma}}}\right)\phi=0;\qquad\rho=\sigma\sinh^{-1}{z},$ (A.2)
where
$\kappa^{2}=E-\frac{1/4+2\eta-3\nu}{\sigma^{2}},\qquad\lambda(\lambda-1)=-(1/4+4\nu-2\eta-\mu^{2}\sigma^{2}).$ (A.3)
For the bound states to exist $\lambda(\lambda-1)>0$. From Eq. (A.3) we find $\lambda=\frac{1}{2}+\sqrt{\mu^{2}\sigma^{2}+2\eta-4\nu}$. Taking into account the solution $\phi$ of Eq. (A.2), we find [33] the general solution (17) in the inner region
[…]
where $A(z)=\left(\frac{\sigma^{2}}{1+z^{2}}\right)^{1/4}(1+z^{2})^{\lambda/2}$, $a=\frac{1}{2}(\lambda+i\kappa\sigma)$, $b=\frac{1}{2}(\lambda-i\kappa\sigma)$. Applying these solutions (A.4) to the transcendental equation (28), we obtain the ‘analytical’ energy spectrum. Fig. (3) displays graphs of both odd and even solutions, while Table (2) presents the energy spectra obtained for various KEOs.
(a) Even wave functions
(b) Odd wave functions
Figure 3: Wave function (17) with $\psi_{\text{in}}^{1}(z)$, $\psi_{\text{in}}^{1}(z)$ given by Eqs. (A.4) for the distributions (A.1). The energy values, $\{E_{0},E_{1},E_{2},E_{3},E_{4},E_{5},E_{6}\}=\{-8.25,-6.875,-5.625,-4.50009,-3.5013,-2.63724,-1.96428\}$, are found through the trascendental equation (28) using the BD-D condition.
On the other hand, we can see that Eq. (A.2) is the Schrödinger equation for a constant mass particle subject to a modified Pöschl-Teller potential. The energy spectrum, $\kappa^{2}$ (i.e. $E:=\mathbf{E}_{n}$), and its solutions [10, 33] in the whole space is well known, that is to say, if $z_{0}\rightarrow$-$\infty$, $z_{1}\rightarrow\infty$ in the model (2) we find
$\mathbf{E}_{n}=-\frac{(n-\lambda+1)^{2}}{\sigma^{2}}+\frac{1/4+2\eta-3\nu}{\sigma^{2}};\qquad n=0,1,2,\dots,\lambda-1$ (A.5)
which corresponds to the case in which the double heterostructure ceases to be so.
A.2 Complement of the DH defined in Table 5.
Let’s examine a second system with analytical solutions, based on the distributions of the DH defined in Table 5
$V_{\textrm{in}}(z)=-V_{0}^{M}e^{\sigma z}(2-e^{\sigma z}),\qquad m_{\textrm{in}}(z)=m_{0}^{M}\sigma^{2}e^{-2\sigma z},$ (A.6)
The associated Schrödinger equation (12) is
$\frac{d^{2}\phi}{d\rho^{2}}+\left(\kappa^{2}-\frac{2V_{0}^{M}\sqrt{m_{0}^{M}}}{\rho}-\frac{3/4+m_{0}^{M}V_{0}^{M}+2\eta-2\nu}{\rho^{2}}\right)\phi=0;\qquad\rho=-\sqrt{m_{0}^{M}}e^{-\sigma z},$ (A.7)
$\kappa^{2}=E$. For the case that concerns us, viz. bound states, E=|E|E=-|E|. Eq. (A.7) can be rewritten as
$\frac{d^{2}\phi}{dy^{2}}+\left(-\frac{1}{4}+\frac{q}{y}+\frac{1/4-r^{2}}{y^{2}}\right)\phi=0,$ (A.8)
where $y=-2i\kappa\rho$, $q=-V_{0}^{M}\sqrt{m_{0}^{M}}/\sqrt{|E|}$, $r^{2}=1+m_{0}^{M}V_{0}^{M}+2\eta-2\nu$. Taken into account the solutions [1], $M_{q,r}(y),W_{q,r}(y)$, of (A.8), the exact solutions (11) for this system (A.6) are
$\psi_{\text{in}}^{1}(z)=\left(m_{0}^{M}\right)^{1/4}\sqrt{\sigma}\,e^{-\sigma z/2}M_{q,r}(y);\qquad\psi_{\text{in}}^{2}(z)=\left(m_{0}^{M}\right)^{1/4}\sqrt{\sigma}\,e^{-\sigma z/2}W_{q,r}(y).$ (A.9)
$M_{q,r}(y)$ and $W_{q,r}(y)$ are the Whittaker functions of first and second kind respectively. Figure (4) illustrates three functions along with their energies calculated from Eq. (28). Table (5) compares these results with values derived from the coefficient poles.
Figure 4: Wave function (17) with $\psi_{\text{in}}^{1}(z)$, $\psi_{\text{in}}^{1}(z)$ given by Eqs. (A.9). The energy values $\{E_{0},E_{1},E_{2}\}$ are found in Table 5.
Appendix B
Another potential solution to Equation (9) is by substituting $\psi(z)=m(z)^{1/2}\Phi$. This substitution leads to the following equation
$\left(-\frac{d^{2}}{dz^{2}}+V^{*}(z)-E^{*}\right)\Phi=0.$ (B.1)
where
$V^{*}(z)-E^{*}:=\frac{1}{4}\frac{m^{\prime}(z)^{2}}{m(z)^{2}}(3+2\eta)-\frac{1}{2}\frac{m^{\prime\prime}(z)}{m(z)}(1+\nu)+(V(z)-E)m(z).$ (B.2)
Eq. (B.1) can be solved as long as $V^{*}(z)$ is an exact, quasi-exact, or conditionally exact potential, and $E^{*}$ is a constant.
B.1 Application to the DH defined in Table 8
For the profiles of DH of Section 4.4
$V_{\textrm{in}}(z)=V_{c}e^{cz};\qquad m_{\textrm{in}}(z)=\mu_{0}e^{cz};\qquad\mu_{0},V_{c}>0.$
we get the equation
$\left(-\frac{d^{2}}{dz^{2}}+2\mu_{0}e^{cz}(V_{c}e^{cz}-E)-E^{*}\right)\Phi=0;\qquad E^{*}=\frac{c^{2}(-1-2\eta+2\nu)}{4}.$ (B.3)
Or
$\left(-\frac{d^{2}}{dx^{2}}+De^{cx}(e^{cx}-2)-E^{*}\right)\Phi=0;\qquad D=\frac{E^{2}\mu_{0}}{V_{c}},\qquad e^{cx}=\frac{2V_{c}}{E}e^{cz}.$ (B.4)
Following the standard procedure, the ansatz $\Phi=y^{s}e^{-y/2}F(y)$ let us to confluent hypergeometric equation
$y\frac{d^{2}F}{dy^{2}}+(b-y)\frac{dF}{dy}-aF=0;\qquad a=\frac{1}{2}+s-\lambda,\qquad b=1+2s,$ (B.5)
where has been defined $y=2\lambda e^{cx}$, $\lambda=\sqrt{D}/c$, $s^{2}=-E^{*}/c^{2}$, and whose solutions are the Kummer’s function of the first kind $M(a,b;y)={}_{1}F_{1}(a,b;y)$ and the Tricomi confluent hypergeometric function
$U(a,b;y)=\frac{\Gamma(1-b)}{\Gamma(a+1-b)}M(a,b;y)+\frac{\Gamma(b-1)}{\Gamma(a)}z^{1-b}M(a+1-b,2-b;y).$ (B.6)
B.2 Application to the DH defined in Table 9
On the other hand, if we used the potencial and mass given by (55)
$m_{\textrm{in}}(z)=cz^{2};\qquad V_{\textrm{in}}(z)=\frac{A}{c}\left(\frac{1}{z^{4}}+\frac{B}{z^{2}}\right),\qquad z,c,A,-B>0.$
we obtain the Schrödinger equation
$-\frac{d^{2}}{dz^{2}}\Phi(z)+\left(\frac{1}{4}\omega^{2}z^{2}+\frac{g}{2z^{2}}-E^{*}\right)\Phi(z)=0,$ (B.7)
whose energy spectra is [4, 35, Cariñena_2008, 18, 30]
$E^{*}=\omega\left(2n+1+\frac{1}{2}\sqrt{1+2g}\right).$ (B.8)
For this example $E^{*}=-AB$, $\omega^{2}=-4cE$, and $g=2(2+2\eta-\nu)+2A$.

### 21 May 2026 v2

@@
 Finally, we show how these methods allow us to critically analyze some statements that have been made about the relevance of some kinetic energy operators in some simpler heterostructures.
-Keywords— Double heterostructure, effective position-depend mass, von Ross’s operator, boundary conditions, kinetic energy operator, energy spectra, reflection and transmission coefficients.
+Keywords— Double heterostructure, effective position-depend mass, von Roos’s operator, boundary conditions, kinetic energy operator, energy spectra, reflection and transmission coefficients.
 1 Introduction
 When considering heterostructures in condensed matter physics, the low-energy excitations of charge carriers are described by an effective Schrödinger equation, where the KEO is an ambiguous quantity.
-This ambiguity is due to the fact that the mass operator $m=M_{0}m(z)$, no longer commutes with the momentum operator $\hat{p}=-i\hbar d/dz$, leaving the called von Ross KEO [34]
+This ambiguity is due to the fact that the mass operator $m=M_{0}m(z)$, no longer commutes with the momentum operator $\hat{p}=-i\hbar d/dz$, leaving the called von Roos KEO [34]
 $T_{\textrm{vR}}(\alpha,\gamma)=\frac{1}{4}(m^{\alpha}\hat{p}m^{\beta}\hat{p}m^{\gamma}+m^{\gamma}\hat{p}m^{\beta}\hat{p}m^{\alpha}),$ (1)
@@
 $T_{L}=\frac{1}{6}\left(m^{-1}\hat{p}^{2}+\hat{p}m^{-1}\hat{p}+\hat{p}^{2}m^{-1}\right),$ (4)
-which has the ambiguity defined by conditions $\alpha+\gamma=-\frac{2}{3}$, $\alpha\gamma=0$ and therefore is a particular case of von Ross KEO with $\alpha\neq\gamma$.
+which has the ambiguity defined by conditions $\alpha+\gamma=-\frac{2}{3}$, $\alpha\gamma=0$ and therefore is a particular case of von Roos KEO with $\alpha\neq\gamma$.
 For this KEO, $T_{L}$, the boundary conditions for the wave function $\psi(z)$ of charge carriers at points $z_{k}$ within the heterostructure, are [23]
@@
 (2)), as described in Secs.
-4.4, and 4.5.
+48, and 4.5.
 Given that from our DH models we can return to the GH model, we critically discuss the procedure followed in Ref [6] to discard KEOs.
@@
 $|X|=0.$ (28)
-This is a transcendental equation that has been obtained through the analytical process described and allows us to calculate the energies of the bound states, as seen later in the DHs of Sections 4.1, 4.2, 4.4, and 4.5.
+This is a transcendental equation that has been obtained through the analytical process described and allows us to calculate the energies of the bound states, as seen later in the DHs of Sections 4.1, 4.2, 48, and 4.5.
 The energy spectra of these analytical bound states are compared with the reflection coefficient poles which can be calculated using the multi-step methodology of Section 3.2.
@@
 In principle, both potentials are of completely different nature:
-  • The term $g/\rho^{2}$ is a repulsive (attractive) potential barrier with a positive (negative) singularity at the origin only when its coefficient $g\geq 0$ ($g<0$).
+  • The term $g/\rho^{2}$ is a repulsive (attractive) potential barrier with a positive (negative) singularity at the origin only when its coefficient $g>0$ ($g<0$).
 Nevertheless, this term retains the isochronous property that characterizes the harmonic oscillator when $g\geq-1/2$.
@@
 It is the BD-D and L operators that determine the lowest energy spectra of the double heterostructures defined in Tables 7 and 8, respectively.
-We have used the ordering of the kinetic energy operator that has been proposed in reference [23] and have also shown that von Ross’s KEO includes it.
+We have used the ordering of the kinetic energy operator that has been proposed in reference [23] and have also shown that von Roos’s KEO includes it.
 We have not found double heterostructures with ambiguities that prevent obtaining well-defined energy spectra, contrary to what has been reported in other references.
@@
 Acknowlegments
-RVT acknowledges Secretaría de Ciencia, Humanidades, Tecnología e Innovación (SECIHTI - México) support under the grant FORDECYT-PRONACES/61533/2020.
+RVT and ECO acknowledges Secretaría de Ciencia, Humanidades, Tecnología e Innovación (SECIHTI - México) support under the grant FORDECYT-PRONACES/61533/2020.
 JGR acknowledges IPN project SIP 20260578 and SECIHTI.
@@
 B.1 Application to the DH defined in Table 8
-For the profiles of DH of Section 4.4
+For the profiles of DH of Section 48
 $V_{\textrm{in}}(z)=V_{c}e^{cz};\qquad m_{\textrm{in}}(z)=\mu_{0}e^{cz};\qquad\mu_{0},V_{c}>0.$